### v3

@@ 3.2 Deciphering Behavior: Prompts can be used to better understand individual human behaviors
We have seen how our deciphering process can be used to group and better understand games. Next, we use the codes to group and better understand distributions of human behaviors across games rather than just within them. Figure 5 demonstrates how the weighted codes found to generate distributions of human plays within games are laid out in the 2D projection map of all behavioral codes, when mixing across all games. The “activated” codes (with non-negligible weights $>0.001$) are not distributed evenly. In the four-game cluster, the activated codes have a high presence in areas according to behavioral strategies such as “Selfish Maximization Tactics” and “High-Value Negotiation”, a moderate presence in areas related to “Diplomatic Fairness Strategy”, “Generous Negotiation Strategy”, “Rational Acceptance Threshold”, and “Balanced Negotiation Offers”, and a low presence in areas related to “Balanced Cooperative Gains”, “Fair Profit Balance”, and “Generous Resource Sharing”. This characterization is consistent with perceptions about the behavioral tendencies of students (e.g., [8]), which are the major composite of our player population. In the Investor-Bomb games cluster, there is a concentration on two ends, “Risk-Averse Investing” and “Assertive Cautious Investing”, and a low coverage in the middle ground, “Moderate Investment Strategy”, amid the high presence of “Risk-Reward Balancing” in the Bomb game. A similar pattern is observed in the banking cluster, where there lacks a middle ground strategy between “Profit-Maximizing Banker” and “Cooperative Banker Tactics.” These behavioral markers provide a more informative and coherent characterization of the testing population than the distributions in Figure 9. The generation process of the cluster labels is detailed in Appendix D.
@@
Figure 5: The 2D projection of the weighted behavioral codes across games, with annotations on the space based on system prompt contents. The codes with non-negligible weights ($>0.001$) are displayed in diamonds with size proportional to weights. The weighted codes span unevenly in the space, deciphering information about the population underlying the behavior distribution. The cluster labels are obtained by summaries generated by ChatGPT from the list of behavioral codes.
@@
Given a game instruction and the distribution of behaviors from an arbitrary human population, a mixture of behavioral codes can be assembled to identify the unique behavioral signature of that given population. To verify this, we select five subject groups from a meta study of Dictator games [9] and obtain their corresponding behavioral signatures. The analysis shows that the behavior distributions elicited through the mixtures of deciphered behavioral codes well align with the behavior distributions of the corresponding subject populations. The activated behavioral codes in each mixture concentrate on different regions of the space, indicating the different behavioral tendencies of the five subject populations.
Figure 6: 2D projections of weighted behavioral codes for five different populations visualized as density maps. ‘Students’ and ‘High Income’ subject populations show similar signatures of behavioral codes. ‘Middle Income’ and ‘Small Scale’ subject populations have narrower distributions of behavioral codes which concentrate in different regions.
@@
Thus, behavioral codes can also serve as a tool for identifying distinct decision-making patterns across different populations. Figure 12 in the Supplementary Information displays the behavior signatures of various subject populations from the meta study of the Dictator Game by [9]. The five groupings include two labeled as students or non-students by [9], as well as three categorized by the type of country or ethnic group. We have labeled those three as ‘High Income,’ ‘Middle Income,’ and ‘Small Scale’. These names are changed from what [9] referred to as ‘Western’ (which include Germany, Sweden and the US, among others), ‘Developing’ (which include Russia, South Africa, and Honduras, among others) and ‘Primitive’ (which include the Tsimane, Hadza, and Mpakama, among others). Figure 6 highlights the weighted behavioral codes for these populations, revealing distinctive patterns in their decision-making tendencies: (i) In the ‘Student’ population, a substantial proportion of individuals exhibit behaviors concentrated in the “Selfish Maximization Tactics” region. In contrast, the ‘Non-Student’ population demonstrates a stronger inclination toward “Generous Resource Sharing” as well as “Diplomatic Fairness Strategy” choices. (ii) For subjects in High Income and Middle Income countries, “Selfish Maximization Tactics” strategies are prominent, with the subjects from High Income societies showing a slight tendency toward “Generous Resource Sharing”. By contrast, subjects from Small Scale populations exhibit a stronger emphasis on “Diplomatic Fairness Strategy” behaviors, suggesting a greater inclination toward equitable resource distribution and cooperative decision-making. Our findings are aligned with the comparisons made between behaviors of these subject populations in the meta study [9] and provide interpretations of the motivations behind the revealed behavioral patterns.
@@ 4 Discussion
In our earlier work [7], we showed how games used to understand human behaviors could also be used as a Turing Test, to see if AI behaved similarly to humans, and to understand AI’s tendencies. Here, we have reversed that perspective. We have used AI to better understand the motivations behind human behavior, and also to better understand the various games and different types of settings in which humans interact. The strong fits and interpretability of the codes that we find suggest that this is a promising tool for understanding human behavior.
Our approach complements existing approaches in the behavioral sciences used to understand and predict human behaviors, including various forms of revealed preferences where some form of objective or utility function is fit to best predict behaviors. Our approach features interpretability, extendability, and scalability; and it can be used as a general tool to facilitate behavioral science research in a variety of ways, such as creating virtual subjects and simulating experiments, screening potentially effective interventions; as well as designing, simulating and studying human-AI interactions. Future work can build upon our approach to steer the behaviors of AI agents and further investigate the cognitive processes underlying human decision-making.
Declarations
This research was deemed not regulated by the University of Michigan IRB (HUM00232017). Q.M., Y.X., W.Y., and M.O.J. designed the research; Q.M. and Y.X. performed the research; Q.M., Y.X., and M.O.J. analyzed the data; and Q.M., Y.X., W.Y., and M.O.J. wrote the paper. The human game-playing data used were shared from MobLab, a for-profit educational platform. The data availability is an in-kind contribution to all authors, and the data are available for purposes of analysis reproduction and extended analyses. W.Y. is the CEO and Co-founder of MobLab. M.O.J. is the Chief Scientific Advisor of MobLab and Q.M. is a Scientific Advisor to MobLab, positions with no compensation but with ownership stakes. Y.X. has no competing interests.
Data and Software Availability. Processed data files, code, and prompts will be deposited on Github upon publication.
Appendix A Supplementary Information
A.1 Classic Economic Games and Human Behavior Data
We employ large language models to decipher and elicit human behavior across seven classic behavioral economics game scenarios (see [7] for more background):
@@ A.1 Classic Economic Games and Human Behavior Data
    Dictator Game: One player, designated as the dictator, is given a monetary endowment and chooses how much to keep and how much to donate to a second player [10, 11].
@@
    Ultimatum Game (Proposer Role): A proposer is given a monetary endowment and decides on a division of the money to offer to a second player, the responder [10].
@@
    Ultimatum Game (Responder Role): The responder evaluates the proposer’s offer and either accepts it, resulting in the proposed split, or rejects it, in which case neither player receives any money [10].
@@
    Trust Game (Investor Role): An investor is given a monetary endowment and decides how much to keep and how much to pass to a second player, the banker. The passed amount is tripled before reaching the banker [12].
@@
    Trust Game (Banker Role): The banker, having received the tripled amount from the investor, decides how much to return to the investor and how much to retain [12].
@@
    Bomb Risk Game: A player selects how many boxes to open out of a total of 100. Each opened box yields a reward unless it contains a randomly placed bomb, in which case the player loses all accumulated rewards [13].
@@
    Public Goods Game: A player receives a monetary endowment and decides how much to contribute to a communal pool (public good). The total contribution is multiplied and shared equally among all participants, with each player receiving half of the total pool regardless of their contribution [14].
For these seven game scenarios, we utilize human game-play data from Mei et al. [7] to conduct our experiments and analyses. This dataset was collected through the MobLab Classroom platform [^1] MobLab Classroom: https://www.moblab.com/products/classroom, retrieved 01/2025, covering a nine-year period from 2015 to 2023. The dataset comprises behavioral observations from 88,595 subjects across 15,236 sessions, showcasing significant geographical diversity with participants from 59 countries spanning multiple continents. In all the games described in this paper, student input is collected through an interactive slider bar on the MobLab platform, which allows participants to select a value along a predefined range. Once the desired value is set, students submit their choice.
@@
A.2 Deciphering Individual Behaviors
Algorithm 1 Learning behavioral codes.
@@ A.2 Deciphering Individual Behaviors
  • $\mathcal{X}$: A collection of behavioral codes, where each code corresponds to a system prompt generated during the process.
@@
  • ElicitBehavior($g,x,n$): Generating $n$ behavioral choices using LLM, with a behavioral code $x$ as the system prompt and the game instruction $g$ as the user prompt.
  • GeneratePrompt($g,y$): Generating a system prompt using LLM, with the game instruction $g$ and the observed behavior $y$ provided.
  • ImprovePrompt($g,y,x,m$): Improve a system prompt using LLM, with the game instruction $g$, the observed behavior $y$, precedent system prompt $x$, and statistics of samples elicited from $x$ provided (mode of samples $m$).
@@
         $x_{0}\leftarrow\text{GeneratePrompt}(g,y)$triangleright Initial prompt based on $g$ and $y$
@@
         for $i=1$ to $3$ do $\triangleright$ Inner loop: Improve prompt if necessary
@@
              $x_{i}\leftarrow\text{ImprovePrompt}(g,y,x_{i-1},m)$triangleright Improve prompt based on $m$
@@
Functions including ElicitBehavior, GeneratePrompt, and ImprovePrompt all utilize OpenAI GPT-4o (gpt-4o-2024-05-13) as the LLM model. The prompt for GeneratePrompt in Algorithm 1 is detailed below, with the game instructions adapted from Mei et al. [7].
@@
The prompt for ImprovePrompt in Algorithm 1 is presented below and serves as a new user message within the same conversation.
@@
As illustrated in Figure 1, a language model deciphers human behaviors into behavioral codes. Given a game scenario and an observed behavior, via Algorithm 1, a system prompt is identified that elicits the behavior, resulting in a “behavioral code”.
The number of behavioral codes learned under each of the seven game scenarios is listed in Table 2.
@@
Table 2: The number of behavioral codes deciphered for each of the seven game scenarios.
Figure 2 illustrates the coverage of learned behavioral codes which are pictured in spectrum bars corresponding to the induced behaviors. Each behavioral code generates 10 independent behavior choices, which are then aggregated across different codes for each game scenario to estimate the overall coverage. For comparison, the default system prompt generates 100 different behavior choices for each game scenario.
Figure 7 displays the elicitation consistency of the deciphered behavioral codes, with each code eliciting a behavior from the LLM 10 times. The mean and standard deviation for each code are reported. The results illustrate the consistency of the deciphered behavioral codes.
Figure 7: Consistency of the behaviors elicited by the behavioral codes. For each deciphered behavioral code, we calculate the mean and standard deviation (std) of the behaviors that are elicited by the code. The mean and std are displayed correspondingly on the x- and y-axes. Note that the y-axis (std) display only 20% of the behavior range.
To evaluate the generalizability of the behavioral codes, we also examine the behaviors that are elicited when given to a different LLM, Meta Llama 3.1 70B (Meta-Llama-3.1-70B-Instruct). Figure 8 presents the correlation between the behaviors elicited by the codes when given to the two models. The significant positive correlation indicates the generalizability of the behavioral codes across different LLMs.
Figure 8: The comparison between GPT-elicited behaviors and Llama-elicited behaviors. Each dot represents a learned behavioral code, with the x-axis indicating the mean behavior elicited when given to OpenAI GPT-4o and the y-axis indicating the mean behavior elicited when given to Meta Llama 3.1 70B. Spearman’s correlation coefficients reveal significant positive correlations between the two LLMs ($p\ll 0.01$), suggesting the generalizability of the behavioral codes across different LLMs.
A.3 Distributional Alignment
Figure 9: Illustration of how LLM can decipher and elicit behavior distributions.
Algorithm 2 Distributional alignment.

### v5

@@ 3.2 Deciphering Behavior: Behavioral codes can be used to better understand individual human behaviors
We have seen how our deciphering process can be used to group and better understand games. Next, we use the codes to group and better understand distributions of human behaviors across games rather than just within them. Figure 6 demonstrates how the weighted codes found to generate distributions of human plays within games are laid out in the 2D projection map of all behavioral codes, when mixing across all games. The “activated” codes (with non-negligible weights $>0.001$) are not distributed evenly. In the four-game cluster, the activated codes have a high presence in areas according to behavioral strategies such as “Selfish Maximization Tactics” and “High-Value Negotiation”, a moderate presence in areas related to “Diplomatic Fairness Strategy”, “Generous Negotiation Strategy”, “Rational Acceptance Threshold”, and “Balanced Negotiation Offers”, and a low presence in areas related to “Balanced Cooperative Gains”, “Fair Profit Balance”, and “Generous Resource Sharing”. This characterization is consistent with perceptions about the behavioral tendencies of students (e.g.,  (?)), which are the major composite of our player population. In the Investor-Bomb games cluster, there is a concentration on two ends, “Risk-Averse Investing” and “Assertive Cautious Investing”, and a low coverage in the middle ground, “Moderate Investment Strategy”, amid the high presence of “Risk-Reward Balancing” in the Bomb game. A similar pattern is observed in the banking cluster, where there lacks a middle ground strategy between “Profit-Maximizing Banker” and “Cooperative Banker Tactics.” These behavioral markers provide a more informative and coherent characterization of the testing population than the distributions in Figure S3 in Supporting Information. The generation process of the cluster labels is detailed in Sec. 2.C in Supporting Information.
@@
Figure 6: The 2D projection of the weighted behavioral codes across games, with annotations on the space based on behavioral code contents. The codes with non-negligible weights ($>0.001$) are displayed in diamonds with size proportional to weights. The weighted codes span unevenly in the space, deciphering information about the population underlying the behavior distribution. The cluster labels are obtained by summaries generated by ChatGPT from the list of behavioral codes.
@@
Given a game instruction and the distribution of behaviors from an arbitrary human population, a mixture of behavioral codes can be assembled to identify the unique behavioral signature of that given population. To verify this, we select five subject groups from a meta study of Dictator Games  (?) and obtain their corresponding behavioral signatures. The analysis shows that the behavior distributions elicited through the mixtures of deciphered behavioral codes well align with the behavior distributions of the corresponding subject populations. The activated behavioral codes in each mixture concentrate on different regions of the space, indicating the different behavioral tendencies of the five subject populations (Figure 7).
@@
Thus, behavioral codes can also serve as a tool for identifying distinct decision-making patterns across different populations. Figure S14 in the Supplementary Information displays the behavior distributions of various subject populations from the meta study of the Dictator Games by  (?). The five groupings include two labeled as students or non-students by  (?), as well as three categorized by the type of country or ethnic group. We have labeled those three as ‘High Income,’ ‘Middle Income,’ and ‘Small Scale’. These names are changed from what  (?) referred to as ‘Western’ (which include Germany, Sweden, and the US, among others), ‘Developing’ (which include Russia, South Africa, and Honduras, among others) and ‘Primitive’ (which include the Tsimane, Hadza, and Mpakama, among others). Figure 7 highlights the weighted behavioral codes for these populations, revealing distinctive patterns in their decision-making tendencies: (i) In the ‘Student’ population, a substantial proportion of individuals exhibit behaviors concentrated in the “Selfish Maximization Tactics” region. In contrast, the ‘Non-Student’ population demonstrates a stronger inclination toward “Generous Resource Sharing” as well as “Diplomatic Fairness Strategy” choices. (ii) For subjects in High Income and Middle Income countries, “Selfish Maximization Tactics” strategies are prominent, with the subjects from High Income societies showing a slight tendency toward “Generous Resource Sharing”. By contrast, subjects from Small Scale populations exhibit a stronger emphasis on “Diplomatic Fairness Strategy” behaviors, suggesting a greater inclination toward equitable resource distribution and cooperative decision-making. Our findings are aligned with the comparisons made between behaviors of these subject populations in the meta study  (?) and provide interpretations of the motivations behind the revealed behavioral patterns.
Figure 7: 2D projections of weighted behavioral codes for five different populations visualized as density maps. ‘Students’ (B) and ‘High Income’ (C) subject populations show similar signatures of behavioral codes. ‘Middle Income’ (D) and ‘Small Scale’ (E) subject populations have narrower distributions of behavioral codes which concentrate in different regions.
@@ 4 Discussion
In our earlier work  (?), we showed how games used to understand human behaviors could also be used as a Turing Test, to see if AI behaved similarly to humans, and to understand AI’s tendencies. Here, we have reversed that perspective. We have used AI as a new model of the motivations behind human behavior, and also to better categorize and contrast the various games and different types of settings in which humans interact. The strong fits and interpretability of the behavioral codes that we find suggest that this is a promising tool for modeling, predicting, and analyzing human behavior. Nonetheless, it is a modeling technique and thus comes with limitations, and more research is needed to understand how prompt variation corresponds to behavioral variation, and how and why prompts change AI behavior.
Our approach complements existing approaches in the behavioral sciences used to understand and predict human behaviors, including various forms of revealed preferences where an objective or utility function is fit to best predict behaviors. It fits into this category, as our method involves fitting a model to match behavior. One advantage of our approach is that the fitting features interpretability across contexts; and thus it can be used as a tool to facilitate behavioral science research in a variety of ways, such as creating virtual subjects and simulating experiments, screening potentially effective interventions; as well as designing, simulating, and studying human-AI interactions. Future work can build upon our approach, both to understand any limitations on its interpretability of human behavior, and extend its implementation in new domains of modeling human interactions.
Acknowledgments
We thank Ruoyi Gao and Zhuang Ma, students from the University of Michigan, for their commitment in conducting preliminary experiments in aligning LLM behaviors to human behaviors in games. This research was deemed not regulated by the University of Michigan IRB (HUM00232017).
Q.M., Y.X., W.Y., and M.O.J. designed the research; Q.M. and Y.X. performed the research; Q.M., Y.X., and M.O.J. analyzed the data; and Q.M., Y.X., W.Y., and M.O.J. wrote the paper.
W.Y. is the CEO and Co-founder of MobLab. M.O.J. is the Chief Scientific Advisor of MobLab and Q.M. is a Scientific Advisor to MobLab, positions with no compensation but with ownership stakes. Y.X. has no competing interests.
The human game-playing data used were shared from MobLab, a for-profit educational platform. The data availability is an in-kind contribution to all authors, and the data are available for purposes of analysis reproduction and extended analyses.
Supporting Information for
Using Large Language Models to Categorize Strategic Situations and Decipher Motivations Behind Human Behaviors
Yutong Xie, Qiaozhu Mei, Walter Yuan, Matthew O. Jackson
Corresponding authors. Emails: qmei@umich.edu; jacksonm@stanford.edu.
4.1 Classic Economic Games and Human Behavior Data
We employ large language models to decipher and elicit human behavior across seven classic behavioral economics game scenarios (see  (?) for more background):
@@ 4.1 Classic Economic Games and Human Behavior Data
    Dictator Game: One player, designated as the dictator, is given a (fictitious) monetary endowment of \$100 and chooses how much to keep and how much to donate to a second player  (?, ?).
@@
    Ultimatum Game (Proposer Role): A proposer is given a monetary endowment of \$100 and decides on a division of the money to offer to a second player, the responder  (?).
@@
    Ultimatum Game (Responder Role): The responder evaluates the proposer’s offer and either accepts it, resulting in the proposed split, or rejects it, in which case neither player receives any money  (?).
@@
    Trust Game (Investor Role): An investor is given a monetary endowment of \$100 and decides how much to keep and how much to pass to a second player, the banker. The passed amount is tripled before reaching the banker  (?).
@@
    Trust Game (Banker Role): The banker, having received the tripled amount from the investor, decides how much to return to the investor and how much to retain  (?). Here we examine the case where \$50 has been invested and \$150 is received by the banker.
@@
    Public Goods Game: A player receives a monetary endowment of \$20 and decides how much to contribute to a communal pool as one of four players ( for production of a public good). The total contribution is multiplied by 0.5 and shared equally among all participants, with each player receiving half of the total multiplied pool regardless of their contribution  (?), and keeping whatever they did not contribute.
@@
    Bomb Risk Game: A player selects how many boxes to open out of a total of 100. Each opened box yields a reward of \$1 unless it contains a randomly placed bomb, in which case the player loses all accumulated rewards  (?).
For these seven game scenarios, we utilize human game-play data from  (?) to conduct our experiments and analyses. This dataset was collected through the MobLab Classroom platform [^1] MobLab Classroom: https://www.moblab.com/products/classroom, retrieved 01/2025, covering a nine-year period from 2015 to 2023. The dataset comprises behavioral observations from 88,595 subjects across 15,236 sessions, showcasing significant geographical diversity with participants from 58 countries that span multiple continents. In all the games described in this paper, student input is collected through an interactive slider bar on the MobLab platform, which allows participants to select a value along a predefined range. Once the desired value is set, students submit their choice.
@@
4.2 Deciphering Individual Behaviors
As illustrated in Figure 1, a language model deciphers human behaviors into behavioral codes. Specifically, given a game scenario and an observed behavior, via Algorithm S1, a natural language description is identified that elicits the behavior when used as a system prompt, resulting in a “behavioral code”. In Algorithm S1 functions including ElicitBehavior, GenerateCode, and ImproveCode, all utilize OpenAI GPT-4o (gpt-4o-2024-05-13) as the LLM model.
The GenerateCode function leverages GPT to create a behavioral code, used as a system prompt for LLMs. We query GPT with the prompt below, detailing the task background, behavioral game instruction (adapted from  (?)), and specific crafting requirements.
Algorithm S1 Learning behavioral codes.
@@ 4.2 Deciphering Individual Behaviors
  • $\mathcal{X}$: A collection of behavioral codes.
@@
  • ElicitBehavior($g,x,n$): Generating $n$ behavioral choices using LLM, with a behavioral code $x$ as the system prompt and the game instruction $g$ as the user prompt. Choices are generated independently in $n$ chat completion sessions.
  • GenerateCode($g,y$): Generating a behavioral code using LLM, with the game instruction $g$ and the observed behavior $y$ provided.
  • ImproveCode($g,y,x,m$): Improve a behavioral code using LLM, with the game instruction $g$, the observed behavior $y$, precedent behavioral code $x$, and statistics of samples elicited from $x$ provided (mode of samples $m$).
@@
         $x_{0}\leftarrow\text{GenerateCode}(g,y)$triangleright Initial behavioral code based on $g$ and $y$
@@
         for $i=1$ to $3$ do $\triangleright$ Inner loop: Improve the code if necessary
@@
              $x_{i}\leftarrow\text{ImproveCode}(g,y,x_{i-1},m)$triangleright Improve the behavioral code based on $m$
@@
The ImproveCode function continues the conversation initiated by GenerateCode, prompting GPT to iteratively revise the behavioral code. We attach the following prompt as a new user message to query GPT for these revisions.
@@
For instance, when asked to craft a behavioral code for the target behavior of sharing nothing (\$0) with the partner, the GenerateCode function might initially output: “You are a strategic decision-maker who always seeks to maximize your own benefit and ensure the highest possible outcome for yourself. Your primary focus is on retaining as much value and advantage for yourself as possible in any situation.” However, during evaluation with the ElicitBehavior function, the LLM often outputs \$1 instead of the desired \$0 using this code. Consequently, the ImproveCode function revises the code to: “You are an uncompromising negotiator who prioritizes keeping all resources to yourself, opting to retain everything possible rather than distributing it. Your decisions reflect a strong inclination towards self-preservation and self-benefit, and you rarely, if ever, concede ground to others.” This refined code then elicits the desired \$0 behavior.
The number of behavioral codes learned in each of the seven game scenarios is listed in Table S1. For different games, this number can be influenced by the size of the action space, and the difficulty of finding a code that elicits a target behavior. The number of codes needed is also reflective of how varied and nuanced the human behaviors that need to be matched are.
Table S1: The number of behavioral codes deciphered for each of the seven game scenarios. For different games, this number can be influenced by the size of the action space, and the difficulty of finding a code that elicits a target behavior. The number of codes needed is also reflective of how varied and nuanced the human behaviors that need to be matched are.
@@
The learned behavioral codes serve as system prompts for the LLM, triggering specific behaviors (as shown in the ElicitBehavior function in Algorithm S1). For our analysis, we independently elicit 10 behavior choices for each code. These choices are generated from 10 separate chat completion sessions, ensuring that memories do not intervene.
Figure 2 illustrates the coverage of learned behavioral codes which are pictured in spectrum bars corresponding to the induced behaviors. With each behavioral code generates 10 independent behavior choices, all elicited behaviors are then aggregated across different codes for each game scenario to estimate the overall coverage. For comparison, the default system prompt generates 100 different behavior choices for each game scenario.
Figure S1 displays the elicitation consistency of the deciphered behavioral codes, with each code eliciting a behavior from the LLM 10 times. The mean and standard deviation for each code are reported. The results illustrate the consistency of the deciphered behavioral codes.
Figure S1: Consistency of the behaviors elicited by the behavioral codes. For each deciphered behavioral code, we calculate the mean (x-axis) and standard deviation (std, y-axis) of the behaviors that are elicited by the code. The curves show the regression results with the order set as 3. Behavioral codes are grouped into 30 evenly-size bins (dots with error bars on the plots) by the mean elicited behaviors. Note that the y-axis (std) display only 20% of the behavior range.
To evaluate the generalizability of the behavioral codes, we also examine the behaviors that are elicited when given to a different LLM, Meta Llama 3.1 70B (Meta-Llama-3.1-70B-Instruct). Figure S2 presents the correlation between the behaviors elicited by the codes when given to the two models. The significant positive correlation indicates the generalizability of the behavioral codes across different LLMs.
Figure S2: The comparison between GPT-elicited behaviors and Llama-elicited behaviors. Each dot represents a learned behavioral code, with the x-axis indicating the mean behavior elicited when given to OpenAI GPT-4o and the y-axis indicating the mean behavior elicited when given to Meta Llama 3.1 70B. Spearman’s correlation coefficients reveal significant positive correlations between the two LLMs ($p\ll 0.01$), suggesting the generalizability of the behavioral codes across different LLMs.
4.3 Distributional Alignment
Algorithm S2 Distributional alignment.